# Enhancing Diffusion-Based Sampling with Molecular Collective Variables

### v1

@@ 1 Introduction
A central goal of statistical mechanics simulations is to estimate macroscopic thermodynamic properties, such as the energy released in a chemical transformation. These simulations typically employ molecular dynamics (MD) or Markov chain Monte Carlo (MCMC), which propagate an atomic structure forward in time or along a chain using energies from quantum mechanical calculations or their approximations (Marx & Hutter, 2000; Frenkel & Smit, 2023; Jacobs et al., 2025). Over long times, the distribution of states converges to the Boltzmann distribution $\nu(x)\propto\exp(-u(x))$, where the reduced energy $u$ encodes the chemical system, temperature, and other thermodynamic variables. Sampling from this distribution is deceptively difficult, despite having access to its explicit unnormalized form, due to the large number of possible atomic configurations and the high energy barriers separating them. Straightforward MD or MCMC often requires an impractically large number of sequential energy evaluations to capture conformational changes or chemical reactions.
Machine learning (ML)-based samplers attempt to bypass dynamics by reformulating sampling as statistical inference on the Boltzmann distribution, using its log-likelihood to draw independent and identically distributed (i.i.d.) samples. Early work focused on normalizing flows (Noé et al., 2019), while recent approaches employ diffusion-based samplers that steer stochastic differential equations (SDEs) whose trajectories yield samples consistent with the target distribution (Zhang & Chen, 2021; Vargas et al., 2023a; Havens et al., 2025). [^1] More information about neural and diffusion-based samplers can be found in Section A.2. Although training can amortize effort across multiple draws, these methods do not inherently reduce the number of energy evaluations required to obtain accurate equilibrium statistics compared to MD-based approaches (He et al., 2025). Directly applying diffusion samplers faces a well-known pitfall of mode collapse, where training and sampling concentrate on high-probability basins while underrepresenting rare but thermodynamically important states. This issue is not only about exploration but also about statistics, since reliable free energies and ensemble averages depend critically on assigning correct weights to rare configurations, which may be exponentially small. Thus, successful sampling must both reach low-occupancy modes and attribute the correct equilibrium weights to them.
@@
Inspired by enhanced sampling methods, we augment the diffusion-based sampler with a bias in collective variable (CV) space. CVs are low-dimensional functions of atomic coordinates that capture slow, chemically relevant motions and can be chosen from common system characteristics (e.g., bond lengths or torsional angles) or identified with ML (Sidky et al., 2020; Bonati et al., 2023). During the training of a diffusion-based sampler, the sampler iteratively proposes configurations and receives feedback from their energies. An additional online repulsive potential is maintained over the CVs: regions visited more often accumulate higher bias, which raises their effective energy, discouraging repeated visits. This flattens free energy barriers over training progress and mitigates collapse onto dominant modes. At inference, importance weights remove the effect of the bias, ensuring that ensemble estimates match the target Boltzmann distribution. The result is broader exploration, including rare modes, while preserving statistically correct estimates after reweighting. Furthermore, compared to MD-based enhanced sampling, the diffusion process mixes more efficiently within the biased distribution, enabling faster exploration. Our contributions are as follows:
  • We propose the Well-Tempered Adjoint Schrödinger Bridge Sampler (WT-ASBS), which augments the state-of-the-art diffusion-based sampler ASBS (Liu et al., 2025) with a biasing mechanism inspired by well-tempered metadynamics (WTMetaD, Barducci et al., 2008), achieving substantially improved exploration and accurate estimation of free energy differences.
@@ 2.1 Molecular Sampling
Biasing-based enhanced sampling approaches often construct a bias that raises the effective sampling temperature along CV space, introduced by Barducci et al. (2008) as the well-tempered target distribution. We first define the CV marginal $\bar{\nu}(s)$ and the potential of mean force (PMF) $F(s)$ from the unbiased CV marginal (defined up to a constant) [^2] The term PMF for $F(s)$ is used interchangeably with the free energy profile/landscape along the CV. using the Dirac delta function $\delta$:
@@
From Equations˜4, 5 and 7, the well-tempered CV marginal is $\bar{\nu}_{\text{WT}}(s)\propto\exp[-\frac{\beta}{\gamma}F(s)]=[\bar{\nu}(s)]^{1/\gamma}$, which shows that the sampled CV behaves as if at a higher effective temperature $T_{\text{eff}}=\gamma T$ ($\beta_{\text{eff}}=\beta/\gamma$), while the orthogonal conditional remains unchanged: $\nu_{\text{WT}}(x\,|\,\xi(x)=s)=\nu(x\,|\,\xi(x)=s)$. This selective increase in sampling temperature along the CV satisfies Requirement˜3. To construct $V_{\text{WT}}$ on-the-fly during simulation, one can stack Gaussian kernels (Barducci et al., 2008), use a kernel density estimator (Invernizzi & Parrinello, 2020), or obtain the bias variationally (Valsson & Parrinello, 2014; Bonati et al., 2019).
@@
Requirement˜2 is sometimes relaxed so that we generate samples uniformly across modes without weighting. This is a practically relevant setting that emphasizes capturing the diversity of metastable modes rather than reproducing their exact Boltzmann probabilities. Mode exploration is the usual objective of data-driven generative models for molecular conformations (Jing et al., 2022), protein structures (Watson et al., 2023), and crystalline materials (Jiao et al., 2023; Miller et al., 2024), which are trained to approximate the empirical distribution of representative structures extracted from exploratory simulations or experimental data without Boltzmann weights.
@@ 2.2 Diffusion-Based Neural Samplers
Diffusion-based neural samplers learn a control vector field that steers a stochastic process to gradually transform noise into a target distribution, typically specified by an unnormalized density or energy function. Unlike iterative generative methods trained on data from implicitly defined distributions (Sohl-Dickstein et al., 2015; Ho et al., 2020; Lipman et al., 2022; Albergo et al., 2023), neural samplers can generate samples by querying the energy alone. This ability is especially relevant for Boltzmann distributions (Zhang & Chen, 2021; Berner et al., 2022; Vargas et al., 2023a; Havens et al., 2025). More information about neural samplers is provided in Section˜A.2.
Adjoint Schrödinger Bridge Sampler (ASBS, Liu et al., 2025)
@@
where $\sigma_{t}$ is the noise schedule, and $p^{u}$ denotes the law of the SDE in Equation˜9.333While the original ASBS formulation is based on SDEs with general drift $f(X_{t},t)+\sigma_{t}u_{\theta}(X_{t},t)$, here we only consider the driftless base process with $f=0$. ASBS avoids backpropagating through the SDE during optimization by introducing the parametrized corrector $h_{\phi}(X_{1})$ that debiases the effect of endpoint coupling from base distribution to allow the SOC interpretation of Equation˜8, which can be converted to Adjoint Matching (AM) objective for the drift (Havens et al., 2025). The corrector is then learned by optimizing the Corrector Matching (CM) objective:
@@
where $\bar{u}:=\text{{stopgrad}}(u)$. These paired objectives instantiate iterative proportional fitting (IPF, Kullback, 1968) and converge to the SB solution. The expectations use closed-form base conditionals, so no target samples or importance weights are needed, which preserves scalability. Furthermore, the control model supports approximate data-based pretraining (see Appendix˜C).
@@ 3.1 Well-Tempered Adjoint Schrödinger Bridge Sampler
We implement WT-ASBS with a well-tempered bias update through stacking Gaussian kernels (Barducci et al., 2008) to reach the well-tempered target Equation˜6. Let $\mathcal{S}\subset\mathbb{R}^{m}$ denote a compact CV region of interest, where bias potentials $V:\mathcal{S}\to\mathbb{R}$ are defined and constructed, which induces the sampling distribution $\nu_{V}(x)\propto\exp[{-\beta(E(x)+V(\xi(x)))}]$. WT-ASBS presented in Algorithm˜1 is executed via a two-time-scale scheme:
@@ 3.2 Recipe for Efficient Sampling of Molecular Systems
In this work, we consider a single reference configuration that defines the accessible component $\mathcal{A}$, corresponding to the initial conformer used to start MD or MCMC chains. Although Algorithm˜1 does not require data samples for training, as in the original ASBS, the model can be pretrained with available samples to warm-start and improve training efficiency. Accordingly, we perform short MD simulations (local sampling) to generate samples from the mode containing the reference configuration and apply bridge matching pretraining (Appendix˜C) to initialize the control model. During early-stage local conformational exploration without energetic barriers, correlated samplers such as MD or MCMC are more efficient because they move locally on the energy landscape, unlike diffusion samplers that must transform noise to physical samples through hundreds of model evaluations for every sample. Since this is only a pretraining stage, high accuracy is unnecessary, and we can further reduce computational cost by using approximate energy evaluations (e.g., classical force fields), faster sampling settings, or transfer operator surrogates (e.g., Schreiner et al., 2023) to create the pretraining data.
@@
Upon adopting ASBS, the algorithm introduces several additional hyperparameters governing training. The first is the choice of low-dimensional CVs that defines the space over which modes are explored. CVs may be selected either from chemical intuition or from topological changes induced by the underlying transformation, as demonstrated in Section˜4. For more complex systems, CVs can also be parameterized by shallow neural networks and trained from available conformations (Sidky et al., 2020; Bonati et al., 2023). Other hyperparameters follow conventions from prior enhanced sampling studies. However, the improved mode-mixing behavior of diffusion samplers permits the use of smaller $\gamma$ values, i.e., the effective temperature along the CVs can be lower. Similarly, the Gaussian height $h$ must be reduced, since diffusion samplers generate uncorrelated samples and thus deposit bias more frequently.
To satisfy Requirement˜1, we use restraint potentials. The specific combination of restraints is system dependent, but their application can be formalized through the lens of chemical isomerism. Since which types of isomerism are suppressed under physical conditions is well understood, these serve as a natural guide (Canfield et al., 2018). We provide a detailed treatment in Section˜A.3, and here illustrate the idea with chirality inversion in proteins and peptides. The $C_{\alpha}$ atom of each amino acid is a chiral center: while both L- and D-configurations exist, nearly all natural amino acids only adopt the L-configuration. As the energy function does not inherently distinguish mirror-symmetric configurations, we introduce flat-bottom harmonic potentials on improper torsional angles defining chirality, thereby biasing sampling toward the physically relevant L-configurations.
@@
Remark˜3.1 implies that the PMF along the CVs can be recovered directly from the final bias. While samples stored in the buffer may be analyzed to characterize configurations associated with specific regions in CV space, to recover the full weighted conformational distribution, one can integrate Equation˜9 to generate samples $X_{i}$ and assign weights $W_{i}=\exp[\beta V^{\ast}(\xi(X_{i}))]$, yielding the self-normalized estimator in Equation˜1. If desired, these samples can be refined prior to weighting via short MD or MCMC using $E+V^{\ast}\circ\xi$ as the energy.
@@ 4 Experiments
where $\mathcal{S}_{i}$ is the CV region of state $i$. We report free energy differences $\Delta F_{ij}=F_{j}-F_{i}$, thereby removing the additive constant as it is the same for all $\mathcal{S}_{i}$. We compute multi-state errors with a common offset. Although weighted sampling required by Requirement˜3 is best assessed via PMFs and free energies, we additionally compare with diffusion samplers on alanine dipeptide in Appendix˜E using standard distributional metrics. Our main comparisons are with ASBS and WTMetaD.
@@ 4.1 Conformational Sampling for Small Peptides
We first benchmark the WT-ASBS method on two widely studied peptide systems, alanine dipeptide (Ala2; Ace–Ala–Nme) and alanine tetrapeptide (Ala4; Ace–Ala3–Nme), which serve as standard testbeds for evaluating molecular sampling methods. Both peptides are modeled with a classical force field with implicit water solvation, with the chirality restraint applied to all Cα atoms (see Appendix˜D). In this work, we focus on sampling the full Cartesian coordinates of the molecules. Although peptides and proteins adopt a broad range of conformations, these arise largely from the combinatorial organization of local backbone torsional states, specifically the $\phi$ and $\psi$ angles (Figure˜2a), as classically illustrated by Ramachandran et al. (1963). Building on this insight, we demonstrate that energy-based neural samplers can be made effective for sampling peptides directly in Cartesian space when augmented with biasing along these torsional angles.
Figure 2: Sampling alanine dipeptide. (a) Molecular structure and torsional CVs ($\phi$ and $\psi$). (b) Pretraining data from a short MD simulation. (c) PMF evolution reconstructed from bias during training. (d) PMF from reweighting. (e) Reference PMF. (f) Convergence of $\Delta F$ over training.
@@
We employ $\phi$ and $\psi$ torsion angles as CVs (Figure˜2a). The conformational landscape is broadly divided into two states along $\phi$ (Figure˜2b). Local sampling within the more populated state is efficient due to the absence of a significant barrier; thus, we easily sample pretraining data from a short MD simulation in this region. The pretrained model produces samples restricted to this state. However, once WT-ASBS training begins, the bias deposition progressively shifts the sampling distribution away from the explored region, enabling the discovery of conformations in the less populated state. Over training progress, the deposition asymptotically converges to the bias potential required to reproduce the PMF (Figure˜2c). Figure˜2d shows the PMF obtained from the generated samples of the fully trained model using weights from the final bias. Both the bias-derived PMF and the reweighted-sample PMF agree closely with the reference PMF (Figure˜2e), whereas the baseline ASBS fails to reproduce the correct distribution in the less populated region (Figure˜7). Moreover, compared to the baseline enhanced sampling method (WTMetaD) with the same bias factor, WT-ASBS achieves convergence of the $\Delta F$ between the two states more accurately.
@@
As seen in the Ala2 case, the main energetic barriers separating conformational states arise from backbone $\phi$ torsions. We therefore employ three CVs, $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$. A long 100-µs reference MD simulation shows that the system occupies eight distinct modes, corresponding to the $2^{3}$ combinations of the three $\phi$ torsional states (Figure˜3b). We initialize from a single conformation in the indicated mode and generate pretraining data from a short MD simulation confined in the mode (Figure˜8). As shown in Figure˜3c, WT-ASBS explores conformational states far more efficiently than the baseline WTMetaD, discovering all eight modes within the early stages of training. This advantage arises from its ability to generate uncorrelated samples from the biased potential, enabling exploration of multiple directions of conformational change simultaneously, whereas MD-based enhanced sampling explores states sequentially in simulation time. Consequently, WT-ASBS also achieves effective convergence of free energies (Figure˜3d), as measured by the mean absolute error (MAE) of free energies up to a constant $\min_{c}\frac{1}{8}\sum_{i}|F_{i}-F_{i}^{\text{ref}}+c|$, within chemical accuracy (1 kcal/mol), demonstrating the practical applicability of WT-ASBS.

### v2

@@ 1 Introduction
A central goal of statistical mechanics simulations is to estimate macroscopic thermodynamic properties, such as the energy released in a chemical transformation. These simulations typically employ molecular dynamics (MD) or Markov chain Monte Carlo (MCMC), which propagate an atomic structure forward in time or along a chain using energies from quantum mechanical calculations or their approximations (marx2000ab; frenkel2023understanding; jacobs2025practical). Over long times, the distribution of states converges to the Boltzmann distribution $\nu(x)\propto\exp(-u(x))$, where the reduced energy $u$ encodes the chemical system, temperature, and other thermodynamic variables. Sampling from this distribution is deceptively difficult, despite having access to its explicit unnormalized form, due to the large number of possible atomic configurations and the high energy barriers separating them. Straightforward MD or MCMC often requires an impractically large number of sequential energy evaluations to capture conformational changes or chemical reactions.
Machine learning (ML)-based samplers attempt to bypass dynamics by reformulating sampling as statistical inference on the Boltzmann distribution, using its log-likelihood to draw independent and identically distributed (i.i.d.) samples. Generally, models that approximate thermodynamic ensembles are known as Boltzmann Generators. The first work in this area applied normalizing flows (noe2019boltzmann), while recent approaches employ diffusion-based samplers that steer stochastic differential equations (SDEs) whose trajectories yield samples consistent with the target distribution (zhang2021path; vargas2023denoising; havens2025adjoint). [^1] More information about neural and diffusion-based samplers can be found in Section A.2. Although training can amortize effort across multiple draws, these methods do not inherently reduce the number of energy evaluations required to obtain accurate equilibrium statistics compared to MD-based approaches (he2025trick). Directly applying diffusion samplers faces a well-known pitfall of mode collapse, where training and sampling concentrate on high-probability basins while underrepresenting rare but thermodynamically important states. This issue is not only about exploration but also about statistics, since reliable free energies and ensemble averages depend critically on assigning correct weights to rare configurations, which may be exponentially small. Thus, successful sampling must both reach low-occupancy modes and attribute the correct equilibrium weights to them.
@@
Inspired by enhanced sampling methods, we augment the diffusion-based sampler with a bias in collective variable (CV) space. CVs are low-dimensional functions of atomic coordinates that capture slow, chemically relevant motions and can be chosen from common system characteristics (e.g., bond lengths or torsional angles) or identified with ML (sidky2020machine; bonati2023unified). During the training of a diffusion-based sampler, the sampler iteratively proposes configurations and receives feedback from their energies. An additional online repulsive potential is maintained over the CVs: regions visited more often accumulate higher bias, which raises their effective energy, discouraging repeated visits. This flattens free energy barriers over training progress and mitigates collapse onto dominant modes. At inference, importance weights remove the effect of the bias, ensuring that ensemble estimates match the target Boltzmann distribution. The result is broader exploration, including rare modes, while preserving statistically correct estimates after reweighting. Furthermore, compared to MD-based enhanced sampling, the diffusion process mixes more efficiently within the biased distribution, enabling faster exploration. Our contributions are as follows:
  • We propose the Well-Tempered Adjoint Schrödinger Bridge Sampler (WT-ASBS), which augments the state-of-the-art diffusion-based sampler ASBS (liu2025adjoint) with a biasing mechanism inspired by well-tempered metadynamics (WTMetaD, barducci2008well), achieving substantially improved exploration and accurate estimation of free energy differences.
@@ 2.1 Molecular Sampling
Biasing-based enhanced sampling approaches often construct a bias that raises the effective sampling temperature along CV space, introduced by barducci2008well as the well-tempered target distribution. We first define the CV marginal $\bar{\nu}(s)$ and the potential of mean force (PMF) $F(s)$ from the unbiased CV marginal (defined up to a constant) [^2] The term PMF for $F(s)$ is used interchangeably with the free energy profile/landscape along the CV. using the Dirac delta function $\delta$:
@@
From Equations˜4, 5 and 7, the well-tempered CV marginal is $\bar{\nu}_{\text{WT}}(s)\propto\exp[-\frac{\beta}{\gamma}F(s)]=[\bar{\nu}(s)]^{1/\gamma}$, which shows that the sampled CV behaves as if at a higher effective temperature $T_{\text{eff}}=\gamma T$ ($\beta_{\text{eff}}=\beta/\gamma$), while the orthogonal conditional remains unchanged: $\nu_{\text{WT}}(x\,|\,\xi(x)=s)=\nu(x\,|\,\xi(x)=s)$. This selective increase in sampling temperature along the CV satisfies Requirement˜3. To construct $V_{\text{WT}}$ on-the-fly during simulation, one can stack Gaussian kernels (barducci2008well), use a kernel density estimator (invernizzi2020rethinking), or obtain the bias variationally (valsson2014variational; bonati2019neural). These requirements also extends to a variety of molecular enhanced sampling algorithms, as discussed in Section˜A.1.
@@
Requirement˜2 is sometimes relaxed so that we generate samples uniformly across modes without weighting. This is a practically relevant setting that emphasizes capturing the diversity of metastable modes rather than reproducing their exact Boltzmann probabilities. Mode exploration is the usual objective of data-driven generative models for molecular conformations (jing2022torsional), protein structures (watson2023novo), and crystalline materials (jiao2023crystal; miller2024flowmm), which are trained to approximate the empirical distribution of representative structures extracted from exploratory simulations or experimental data without Boltzmann weights.
@@ 2.2 Diffusion-Based Neural Samplers
Diffusion-based neural samplers learn a control vector field that steers a stochastic process to gradually transform noise into a target distribution, typically specified by an unnormalized density or energy function. Unlike iterative generative methods trained on data from implicitly defined distributions (sohl2015deep; ho2020denoising; lipman2022flow; albergo2023stochastic), neural samplers can generate samples by querying the energy alone. This ability is especially relevant for Boltzmann distributions (zhang2021path; berner2022optimal; vargas2023denoising; havens2025adjoint). More information about neural samplers is provided in Section˜A.2.
Adjoint Schrödinger Bridge Sampler (ASBS, liu2025adjoint)
@@
where $\sigma_{t}$ is the noise schedule, and $p^{u}$ denotes the law of the SDE in Equation˜9.333While the original ASBS formulation is based on SDEs with general drift $f(X_{t},t)+\sigma_{t}u_{\theta}(X_{t},t)$, here we only consider the driftless base process with $f=0$. ASBS avoids backpropagating through the SDE during optimization by introducing the parametrized corrector $h_{\phi}(X_{1})$ that debiases the effect of endpoint coupling from base distribution to allow the SOC interpretation of Equation˜8, which can be converted to Adjoint Matching (AM) objective for the drift (havens2025adjoint). The corrector is then learned by optimizing the Corrector Matching (CM) objective:
@@
where $\bar{u}:=\text{{stopgrad}}(u)$. These paired objectives instantiate iterative proportional fitting (IPF, kullback1968probability) and converge to the SB solution. The expectations use closed-form base conditionals, so no target samples or importance weights are needed, which preserves scalability. Furthermore, the control model supports approximate data-based pretraining (see Appendix˜C).
@@ 3.1 Well-Tempered Adjoint Schrödinger Bridge Sampler
We implement WT-ASBS with a well-tempered bias update through stacking Gaussian kernels (barducci2008well) to reach the well-tempered target Equation˜6. Let $\mathcal{S}\subset\mathbb{R}^{m}$ denote a compact CV region of interest, where bias potentials $V:\mathcal{S}\to\mathbb{R}$ are defined and constructed, which induces the sampling distribution $\nu_{V}(x)\propto\exp[{-\beta(E(x)+V(\xi(x)))}]$. WT-ASBS presented in Algorithm˜1 is executed via a two-time-scale scheme:
@@ 3.2 Recipe for Efficient Sampling of Molecular Systems
In this work, we consider a single reference configuration that defines the accessible component $\mathcal{A}$, corresponding to the initial conformer used to start MD or MCMC chains. Although Algorithm˜1 does not require data samples for training, as in the original ASBS, the model can be pretrained with available samples to warm-start and improve training efficiency. Accordingly, we perform short MD simulations (local sampling) to generate samples from the mode containing the reference configuration and apply bridge matching pretraining (Appendix˜C) to initialize the control model. During early-stage local conformational exploration without energetic barriers, correlated samplers such as MD or MCMC are more efficient because they move locally on the energy landscape, unlike diffusion samplers that must transform noise to physical samples through hundreds of model evaluations for every sample. Since this is only a pretraining stage, high accuracy is unnecessary, and we can further reduce computational cost by using approximate energy evaluations (e.g., classical force fields), faster sampling settings, or transfer operator surrogates (e.g., schreiner2024implicit) to create the pretraining data.
@@
Upon adopting ASBS, the algorithm introduces several additional hyperparameters governing training. The first is the choice of low-dimensional CVs that defines the space over which modes are explored. CVs may be selected either from chemical intuition or from topological changes induced by the underlying transformation, as demonstrated in Section˜4. For more complex systems, CVs can also be parameterized by shallow neural networks and trained from available conformations (sidky2020machine; bonati2023unified). Other hyperparameters follow conventions from prior enhanced sampling studies. However, the improved mode-mixing behavior of diffusion samplers permits the use of smaller $\gamma$ values, i.e., the effective temperature along the CVs can be lower. Similarly, the Gaussian height $h$ must be reduced, since diffusion samplers generate uncorrelated samples and thus deposit bias more frequently.
To satisfy Requirement˜1, we use restraint potentials. The specific combination of restraints is system dependent, but their application can be formalized through the lens of chemical isomerism. Since which types of isomerism are suppressed under physical conditions is well understood, these serve as a natural guide (canfield2018new). We provide a detailed treatment in Section˜A.4, and here illustrate the idea with chirality inversion in proteins and peptides. The $C_{\alpha}$ atom of each amino acid is a chiral center: while both L- and D-configurations exist, nearly all natural amino acids only adopt the L-configuration. As the energy function does not inherently distinguish mirror-symmetric configurations, we introduce flat-bottom harmonic potentials on improper torsional angles defining chirality, thereby biasing sampling toward the physically relevant L-configurations.
@@
Remark˜3.1 implies that the PMF along the CVs can be recovered directly from the final bias. While samples stored in the buffer may be analyzed to characterize configurations associated with specific regions in CV space, to recover the full weighted conformational distribution, one can integrate Equation˜9 to generate samples $X_{i}$ and assign weights $W_{i}=\exp[\beta V^{\ast}(\xi(X_{i}))]$, yielding the self-normalized estimator in Equation˜1. Note that this is exact only in the limit of a perfectly trained sampler. In principle, the path importance weights in ASBS could be used to debias an imperfect sampler with the gradient-based parametrization of control and corrector (liu2025adjoint), though these weights may suffer from high variance. The learned bias also enables a practical refinement strategy: short MD or MCMC trajectories are initiated from the generated samples under the biased energy $E+V^{\ast}\circ\xi$, and the resulting configurations are reweighted with $V^{\ast}$. This restores asymptotic correctness while keeping the procedure computationally efficient thanks to the smoother energy landscape along the CVs.
@@ 4 Experiments
where $\mathcal{S}_{i}$ is the CV region of state $i$. We report free energy differences $\Delta F_{ij}=F_{j}-F_{i}$, thereby removing the additive constant as it is the same for all $\mathcal{S}_{i}$. We compute multi-state errors with a common offset. Although weighted sampling required by Requirement˜3 is best assessed via PMFs and free energies, we additionally compare with diffusion samplers on alanine dipeptide in Appendix˜F using standard distributional metrics. Our main comparisons are with ASBS and WTMetaD.
@@ 4.1 Conformational Sampling for Small Peptides
We first benchmark the WT-ASBS method on two widely studied peptide systems, alanine dipeptide (Ala2; Ace–Ala–Nme) and alanine tetrapeptide (Ala4; Ace–Ala3–Nme), which serve as standard testbeds for evaluating molecular sampling methods. Both peptides are modeled with a classical force field with implicit water solvation, with the chirality restraint applied to all Cα atoms (see Appendix˜D). In this work, we focus on sampling the full Cartesian coordinates of the molecules. Although peptides and proteins adopt a broad range of conformations, these arise largely from the combinatorial organization of local backbone torsional states, specifically the $\phi$ and $\psi$ angles (Figure˜2a), as classically illustrated by ramachandran1963stereochemistry. Building on this insight, we demonstrate that energy-based neural samplers can be made effective for sampling peptides directly in Cartesian space when augmented with biasing along these torsional angles.
Figure 2: Sampling alanine dipeptide. (a) Molecular structure of alanine dipeptide and definition of the torsional collective variables $\phi$ and $\psi$. (b) Distribution of $\phi$ and $\psi$ obtained from a short unbiased MD trajectory used for pretraining the sampler. (c) Evolution of the PMF reconstructed from the accumulated bias during training, illustrating progressive recovery of the major basins in the free energy landscape. The dotted box delineates the two states, one with $\phi<0$ or $\phi>2$ and the other with $\phi\in[0,2]$. (d) PMF estimated by reweighting $10^{6}$ samples generated by the trained sampler using the final bias. (e) Reference PMF computed from a long unbiased MD simulation for comparison. (f) Convergence of free energy differences $\Delta F$ between two states as a function of number of energy evaluations during training.
@@
We employ $\phi$ and $\psi$ torsion angles as CVs (Figure˜2a). The conformational landscape is broadly divided into two states along $\phi$ (Figure˜2b). Local sampling within the more populated state is efficient due to the absence of a significant barrier; thus, we easily sample pretraining data from a short MD simulation in this region. The pretrained model produces samples restricted to this state. However, once WT-ASBS training begins, the bias deposition progressively shifts the sampling distribution away from the explored region, enabling the discovery of conformations in the less populated state. Over training progress, the deposition asymptotically converges to the bias potential required to reproduce the PMF (Figure˜2c). Figure˜2d shows the PMF obtained from the generated samples of the fully trained model using weights from the final bias. Both the bias-derived PMF and the reweighted-sample PMF agree closely with the reference PMF (Figure˜2e), whereas the baseline ASBS fails to reproduce the correct distribution in the less populated region (Figure˜14). Moreover, compared to the baseline enhanced sampling method (WTMetaD) with the same bias factor, WT-ASBS achieves convergence of the $\Delta F$ between the two states more accurately.
To quantify the robustness of WT-ASBS and clarify practical choices, we performed ablations of the WT bias hyperparameters and training mechanisms, and reported detailed results in Appendix˜E. Varying the Gaussian height $h$, kernel width $\sigma$, and bias factor $\gamma$ over wide ranges leaves the final $\Delta F$ unchanged except at the most extreme $h$, while larger $h$ and broader $\sigma$ mainly accelerate early exploration and barrier crossing. The training mechanism ablations show that overly frequent replay buffer refreshes slow convergence because the sampler does not have time to adapt to the evolving bias, whereas moderate update intervals and localized pretraining lead to a faster approach to the reference $\Delta F$ and reduce the total number of energy evaluations. Furthermore, Section˜F.4 demonstrates that WT-ASBS can also operate with an automatically learned ML CV (trizio2021enhanced), recovering accurate PMF and $\Delta F$ after reweighting, which indicates that the method could still be utilized when suitable low-dimensional CVs are not known a priori.
@@
As seen in the Ala2 case, the main energetic barriers separating conformational states arise from backbone $\phi$ torsions. We therefore employ three CVs, $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$. A long 100-µs reference MD simulation shows that the system occupies eight distinct modes, corresponding to the $2^{3}$ combinations of the three $\phi$ torsional states (Figure˜3b). We initialize from a single conformation in the indicated mode and generate pretraining data from a short MD simulation confined in the mode (Figure˜15). As shown in Figure˜3c, WT-ASBS explores conformational states far more efficiently than the baseline WTMetaD, discovering all eight modes within the early stages of training. This advantage arises from its ability to generate uncorrelated samples from the biased potential, enabling exploration of multiple directions of conformational change simultaneously, whereas MD-based enhanced sampling explores states sequentially in simulation time. Consequently, WT-ASBS also achieves effective convergence of free energies (Figure˜3d), as measured by the mean absolute error (MAE) of free energies up to a constant $\min_{c}\frac{1}{8}\sum_{i}|F_{i}-F_{i}^{\text{ref}}+c|$, within chemical accuracy (1 kcal/mol), demonstrating the practical applicability of WT-ASBS. We note that it does not yield lower free energy MAE than WTMetaD, likely because MD already provides efficient local mixing within each basin once barriers along the chosen CV are crossed, whereas the diffusion sampler must learn the full high-dimensional intra-basin distribution. Measure-transport-based global sampling is most beneficial for crossing large barriers, whereas MD-based local sampling approaches can still perform well when refining physical observables, which further suggests that combining the two strategies may be advantageous for more complex systems.